OpenStudy (anonymous):

[10.03] Square root question.

OpenStudy (angela210793):

\[\sqrt{10.03}?\]

OpenStudy (anonymous):

is that the question? if so use a calculator to compute the decimal approximation because this is not a rational number

OpenStudy (anonymous):

Which expression is the simplified form of \[\sqrt{m}^20\]

OpenStudy (anonymous):

Needs to say square root of m to the 20th power.

OpenStudy (anonymous):

COME ON SOMEBODY! ANSWER THE QUESTION. :(

OpenStudy (anonymous):

I think it's m^10 but I'm not sure.

OpenStudy (anonymous):

\[\sqrt{m^{20}}\] is that the question?

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

ok you have it. half of 20 is 10, so your answer is \[m^{10}\] just like you thought

OpenStudy (anonymous):

I believe the answer is M^10 but I'm not sure.